OpenStudy (anonymous):

Compute the first 6 terms in the following sequence:

OpenStudy (anonymous):

\[c _{n}= n + (n+1) + (n+2) +... + (2n)\]

OpenStudy (anonymous):

start with n=1

OpenStudy (helder_edwin):

for instance if n=7 u gotta go from 7 to 2n=14 so \[ \large c_7=7+8+9+10+11+12+13+14 \]

OpenStudy (anonymous):

so for n=1 i would have 1+2=3

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

what about c2

OpenStudy (anonymous):

omg so over thinking that!! i would have 2+3+4=9

OpenStudy (helder_edwin):

well done